# Measurement-induced phase transition in periodically driven free-fermionic systems

### v1

@@ Abstract
It is well known that continuous monitoring of a free-fermionic system leads to a steady state entanglement transition from a logarithmically entangled critical to the area law phase. While unitary evolution tends to grow the entanglement, continuous monitoring opposes this growth and results in the pinning of the wavefunction trajectories to the eigenstate of the measurement operators. In this work, we study the fate of this measurement-induced phase transition in a periodically driven free-fermionic quantum system, where the hopping amplitude is varied periodically in time. We show, in the high-frequency limit, that a renormalization group analysis of the non-Hermitian quantum sine-Gordon model [as proposed in Phys. Rev. X 11, 041004 (2021)] reveals a gapless critical phase with logarithmic entanglement entropy growth and a gapped area-law phase, separated by a Berezinskii-Kosterlitz-Thouless (BKT) transition, which is in excellent agreement with our numerical results. Further, the numerical evidence suggests that such BKT transition also prevails in the low-frequency regime, and decreasing frequency tends to favor the critical phase. Strikingly, we also find that if the hopping amplitude is varied completely symmetrically around zero in the form of square-pulse and sinusoidal drive, the BKT transition is absent in the thermodynamic limit, and the system always favors the area-law phase, independent of the frequency regime.
@@ I Introduction
Understanding the entanglement properties of quantum many-body systems has become a key focus of research, particularly in relation to different phases of matter and the transitions between them, as well as identifying quantum chaos at the many-body level Vidmar and Rigol (2017); Vidmar et al. (2017, 2018); Hackl et al. (2019); Murthy and Srednicki (2019); LeBlond et al. (2019); Modak and Nag (2020); Eisert et al. (2010); Modak and Mandal (2021). Thanks to recent progress on the direct measurement of entanglement dynamics with cold atoms and trapped ions, a key question of how entanglement spreads in out-of-equilibrium many-body systems brought lots of attention in the last decade Islam et al. (2015); Kaufman et al. (2016). Entanglement entropy, a quantifier of quantum correlations between subsystems, exhibits distinct scaling behaviors that provide insight into the nature of the underlying phase. For example, in gapped quantum system, the ground state entanglement entropy typically obeys an area law (scales with the boundary area of the subsystem); in contrast, in one-dimensional critical systems, where long-range correlations are prevalent, it follows a logarithmic scaling with subsystem size, reflecting the scale-invariant nature of these systems at criticality Osterloh et al. (2002); Calabrese and Cardy (2004); Pollmann et al. (2009). On the other hand, under unitary evolution, usually for clean short-ranged systems, the entanglement growth is algebraic in time before saturating to a value that scales linearly with the subsystem size, following a volume law Alba and Calabrese (2017); Calabrese and Cardy (2004); Chapman et al. (2019); Modak et al. (2020). However, introducing disorder into systems changes the entanglement behavior drastically. In a closed, non-interacting quantum system, even small amounts of disorder—whether in one dimension or two—can prevent the thermalization of single-particle states, leading to a phenomenon known as Anderson localization Anderson (1958). Furthermore, when interactions are introduced into these disordered systems, the many-body localized (MBL) phase emerges Basko et al. (2006); Oganesyan and Huse (2007); Žnidarič et al. (2008); Pal and Huse (2010); Serbyn et al. (2015); Modak and Mukerjee (2015); Gong et al. (2021); Bardarson et al. (2012a). In the MBL phase, entanglement grows logarithmically over time, despite the absence of energy transport Bardarson et al. (2012b); Serbyn et al. (2013); Prasad and Garg (2022), in stark contrast to thermal systems, where entanglement growth is typically linear Kim and Huse (2013). Moreover, the typical eigenstate entanglement entropy behaves quite differently in the thermal (ergodic) phase, which follows volume law scaling, compared to the non-ergodic MBL phase, where the scaling follows an area law Luitz et al. (2015). Studies of the fate of MBL are not limited to closed systems; there are several interesting results reported where the system is coupled to a bath Wu and Eckardt (2019); Lüschen et al. (2017); Levi et al. (2016); Fischer et al. (2016); Medvedyeva et al. (2016); Hamazaki et al. (2019); Ren et al. (2020).
@@
Another class of entanglement phase transitions arises from the continuous monitoring of quantum systems. The stochastic Schrödinger equation typically governs these open quantum dynamics, and averaging over different realizations of the states leads to the well-known Lindblad master equation dynamics Paule (2018). Since entanglement entropy is a nonlinear function of the state, calculating it for each trajectory separately and then averaging reveals a phase transition. In contrast, any linear function of the state is featureless and does not exhibit such transitions. Several studies have explored these entanglement phase transitions in both interacting and non-interacting systems, considering clean as well as disordered systems with long-range and short-range hopping, using both Lindblad jump operators and quantum state diffusion approaches Alberton et al. (2021); Szyniszewski et al. (2023); Cao et al. (2019); Minato et al. (2022a); Müller et al. (2022); Yamamoto and Hamazaki (2023); Fuji and Ashida (2020); Ladewig et al. (2022). Continuously monitored open systems display competition between unitary evolution, which drives entanglement growth, and measurements, which tend to localize the wavefunction by collapsing it into eigenstates of the measurement operators. As a result, these systems undergo a measurement-induced phase transition as a function of measurement strength, which also has been reported recently in experiments Koh et al. (2023). It has been shown that, in both clean and disordered free-fermionic systems, there is an entanglement phase transition from a critical to an area law phase as the strength of continuous measurements is tuned. This transition falls into the Berezinskii-Kosterlitz-Thouless universality class Cao et al. (2019); Szyniszewski et al. (2023); Alberton et al. (2021), a result that is also supported by a theoretical study in which the dynamics of a continuously monitored free fermionic system were described by an effective non-Hermitian sine-Gordon model Buchhold et al. (2021). There are recent theoretical studies of a monitored chain of a system of size $L$ of Majorana fermions, which shows ($\ln L)^{2}$ behavior in the critical phase Fava et al. (2023). However, it is important to point out that the stability of the critical phase in a one-dimension free-fermionic system under random projective measurements has been questioned recently, and it has been argued in the thermodynamic limit, for any finite measurement strength, only the area-law phase will survive Poboiko et al. (2023). On the other hand, in the case of interacting systems, a transition from a volume law to an area law has been observed Fuji and Ashida (2020); Yamamoto and Hamazaki (2023); Xing et al. (2024). Furthermore, the role of measurements in localizing quantum states in nonunitary quantum circuits has also attracted significant interest Zabalo et al. (2022); Sierant and Turkeshi (2022); Oshima and Fuji (2023); Weinstein et al. (2023); Côté and Kourtis (2022); Block et al. (2022); Sang and Hsieh (2021); Lunt et al. (2021); Li et al. (2021); Fan et al. (2021); Zabalo et al. (2020); Turkeshi et al. (2020); Gullans and Huse (2020a, b); Bao et al. (2020); Choi et al. (2020); Tang and Zhu (2020); Jian et al. (2020). Some of these studies have also been extended in dimension $D>1$  Chahine and Buchhold (2024); Jin and Martin (2024) and a semi-classical limit of such transition has also been introduced recently Ruidas and Banerjee (2024). In contrast to disorder-induced localization observed in the Anderson-localized or MBL phase, the measurement-induced localization is governed by a completely different mechanism and driven by quantum measurements and measurement strength.
@@
In parallel, in the last few decades, periodically driven (Floquet) quantum systems have emerged as a powerful framework for studying non-equilibrium phases of matter Banerjee and Sengupta (2024); Sarkar et al. (2021, 2022); Nag et al. (2014); Agarwala et al. (2016); Luitz et al. (2018). Floquet systems are known to exhibit exotic phases, such as discrete-time crystals Khemani et al. (2016); Else et al. (2016); Zhang et al. (2017) and Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011); Kundu et al. (2014); Mukherjee et al. (2016). The fate of many-body localization (MBL) in the presence of a periodic drive has also been studied Lazarides et al. (2015); Ponte et al. (2015). While it has been demonstrated that Floquet systems can stabilize or even induce new phases, it remains an open question how continuous monitoring affects these driven systems and whether the introduction of periodic drive can significantly alter the dynamics of a continuously monitored system, either enhancing or suppressing the effects of measurements. The effect of time-dependent Hamiltonian has been studied within the Lindblad master equation formalism Scopa et al. (2019); Ikeda et al. (2021), but the effect of periodic drive on a continuously monitored open quantum system, particularly in terms of trajectory-averaged entanglement entropy, remains unexplored. In our study, we investigate such cases, and our main finding is that periodically driven non-interacting systems also go through BKT transition. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that the critical measurement strength (transition point) increases with the time period. We also find that the effective dynamics of continuously monitored free-fermionic systems with no drive can mimic the high-frequency limit. Moreover, the symmetric drive of the hopping amplitude about a zero mean is special, such a system always favors the area-law phase in the thermodynamic limit, independent of the frequency regime, in contrast to what has been observed in the case of disorder-induced scenario Sarkar et al. (2021). We summarize our set-up and main results in the form of a schematic in Fig. 1.
@@
This manuscript is structured as follows. In Sec. II, we describe our system and protocols. Then, to build the intuitions about our systems, we focus on a $2\times 2$ toy model in Sec. III. Section. IV is dedicated to the many-body systems, which consists of the analytical prediction for high-frequency limit in sec. IV.1, as well as the numerical results in sec. IV.2. Section. IV.3 concentrates on a particular drive protocol that is symmetric around zero means. Finally, we summarize our results in sec. V.
@@
Figure 1: Schematic of our protocol, where a free-fermionic lattice is continuously monitored, and the hopping amplitude $J(t)$ is periodically varied. Phase diagrams for the critical and area-law phases are shown for different time periods, and $\epsilon/2J=1$ corresponds to no drive result.
@@ II Model and Methodology
where, $\hat{c}^{{\dagger}}_{i}$ and $\hat{c}_{i}$ are real-space fermionic creation and annihilation operators respectively. We study the effect of two different kinds of driving with the time period (frequency) $T$ ($\omega=1/T$): one is square pulse drive, where $J(t)=+J$ $(-J+\epsilon)$ for $20\leq t<T/$ 2 ($T/2\leq t<T$), with $\epsilon$ varying between 00 to $2J$ (see Fig. 1); the other is the sinusoidal drive where $J(t)=J\sin(2\pi t/T)$. Note that in the limit $\epsilon=0$, the mean of $J(t)$ is 00 over a time period. We examine this limit separately in this manuscript. We primarily present the results for the square pulse in the subsequent sections of the manuscript; only at the end, in Sec. IV.3, some results for the sinusoidal drive are presented. For all our numerical calculations, we choose $J=1/2$. The time evolution of the state can be described using the stochastic Schrodinger equation as
@@
The above equation describes the continuous monitoring of the particle number $\hat{n}_{i}=\hat{c}_{i}^{{\dagger}}\hat{c}_{i}$ on each site, with measurement strength $\gamma$ Cao et al. (2019). $dW_{i}$s are taken from uncorrelated normal distributions having mean 00 and variance $\gamma dt$, which implies, $\overline{dW_{i,t}}=0$     and    $\overline{dW_{i,t}dW_{j,t^{\prime}}}=\gamma dt\delta_{i,j}\delta(t-t^{\prime})$. For our model and protocols, the total number of particles $N=\sum_{i}\langle\hat{n}_{i}\rangle$ remain conserved in the entire time evolution process; hence, if we start with an initial state with a particular filling at $t=0$, for any time $t>0$, the filling fraction remains the same. In our work, we have fixed the filling fraction $N/L=1/2$ for all many-body calculations. Also, note that the last two terms in Eq. (LABEL:nlse) describe the contribution from the continuous measurement process and can be realized by homodyne detection in quantum optics Wiseman and Milburn (1993); Øksendal and Øksendal (2003).
@@
where, $\mathcal{M}=\sum_{i}[dW_{i}^{t}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt]\hat{n}_{i}$, and $\mathcal{N}$ is the normalization constant. Any pure Gaussian state can be represented as,
@@
where $H(t)=J(t)\sum_{i=1}^{L}(\hat{c}_{i}^{{\dagger}}\hat{c}_{i+1}+h.c.)$ is the single particle Hamiltonian and $M_{ij}=\delta_{i,j}[dW_{i}^{t}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt]$. Under this non-Hermitian evolution, we need to properly normalize the wavefunction in each step, which can be done using the $QR$ decomposition of the $U$ matrix in each step as $U(t+dt)=QR$, and we set the new $U$ as $Q$. Note that $U$ is an $L\crossproduct N$ matrix, follows $U^{{\dagger}}U=\mathbb{I}_{N\crossproduct N}$. On the other hand, the correlation matrix $D$ can be defined as $D_{ij}=\langle\hat{c}_{i}^{{\dagger}}\hat{c}_{j}\rangle=[UU^{{\dagger}}]_{ij}$ For our many-body calculations, we choose $dW_{i}$s from different random realizations, compute the nonlinear function of states, e.g., entanglement entropy using the correlation matrix for each realization separately, and finally average over different realizations (we ensure the convergence of our results by varying the realization number). On the other hand, the evolution of the average state will be governed by the Lindblad equation Cao et al. (2019).

### v2

@@ Abstract
It is well known that unitary evolution tends to increase entanglement, whereas continuous monitoring counteracts this growth by pinning the wavefunction trajectories to the eigenstates of the measurement operators. In this work, we investigate the fate of the measurement-induced phase transition in a periodically driven free-fermionic quantum system, where the hopping amplitude is modulated periodically in time using a square pulse. In the high-frequency limit, a renormalization group analysis of the non-Hermitian quantum sine-Gordon model [as proposed in Phys. Rev. X 11, 041004 (2021)] reveals that if the hopping amplitude is varied symmetrically around zero, the system always favors the area-law phase, where the steady-state entanglement entropy is independent of subsystem size. In contrast, asymmetry in the drive amplitudes tends to promote entanglement growth. Furthermore, numerical evidence for the system sizes accessible to us suggests that decreasing the drive frequency typically favors entanglement growth. For such driven systems, at least for reasonably small frequency regimes, as a function of measurement strength, we observe a potential signature of a Berezinskii-Kosterlitz-Thouless (BKT) phase transition between a gapless critical phase, characterized by logarithmic growth of entanglement entropy with subsystem size, and a gapped area-law phase. However, it is almost impossible to rule out the possibility that the transition observed here is not an actual thermodynamic transition, but a finite-size crossover between logarithmic to area law entanglement phase. Even in that scenario, the critical length scale beyond which the area law phase prevails increases with the increasing time period of driving. On the other hand, for a symmetric drive, the system consistently exhibits an area-law phase, regardless of the driving frequency.
@@ I Introduction
Understanding the entanglement properties of quantum many-body systems has become a key focus of research, particularly in relation to different phases of matter and the transitions between them, as well as identifying quantum chaos at the many-body level Vidmar and Rigol (2017); Vidmar et al. (2017, 2018); Hackl et al. (2019); Murthy and Srednicki (2019); LeBlond et al. (2019); Modak and Nag (2020); Eisert et al. (2010); Modak and Mandal (2021). Thanks to recent progress on the direct measurement of entanglement dynamics with cold atoms and trapped ions, a key question of how entanglement spreads in out-of-equilibrium many-body systems brought lots of attention in the last decade Islam et al. (2015); Kaufman et al. (2016). Entanglement entropy, a quantifier of quantum correlations between subsystems, exhibits distinct scaling behaviors that provide insight into the nature of the underlying phase. For example, in gapped quantum system, the ground state entanglement entropy typically obeys an area law (scales with the boundary area of the subsystem); in contrast, in one-dimensional critical systems, where long-range correlations are prevalent, it follows a logarithmic scaling with subsystem size, reflecting the scale-invariant nature of these systems at criticality Osterloh et al. (2002); Calabrese and Cardy (2004); Pollmann et al. (2009). On the other hand, under unitary evolution, usually for clean short-ranged systems, the entanglement growth is algebraic in time before saturating to a value that scales linearly with the subsystem size, following a volume law Alba and Calabrese (2017); Calabrese and Cardy (2004); Chapman et al. (2019); Modak et al. (2020). However, introducing disorder into systems changes the entanglement behavior drastically. In a closed, non-interacting quantum system, even small amounts of disorder—whether in one dimension or two—can prevent the thermalization of single-particle states, leading to a phenomenon known as Anderson localization Anderson (1958). Furthermore, when interactions are introduced into these disordered systems, the many-body localized (MBL) phase emerges Basko et al. (2006); Oganesyan and Huse (2007); Žnidarič et al. (2008); Pal and Huse (2010); Serbyn et al. (2015); Modak and Mukerjee (2015); Gong et al. (2021); Bardarson et al. (2012a). In the MBL phase, entanglement grows logarithmically over time, despite the absence of energy transport Bardarson et al. (2012b); Serbyn et al. (2013); Prasad and Garg (2022), in stark contrast to thermal systems, where entanglement growth is typically linear Kim and Huse (2013). Moreover, the typical eigenstate entanglement entropy behaves quite differently in the thermal (ergodic) phase, which follows volume law scaling, compared to the non-ergodic MBL phase, where the scaling follows an area law Luitz et al. (2015a). Studies of the fate of MBL are not limited to closed systems; there are several interesting results reported where the system is coupled to a bath Wu and Eckardt (2019); Lüschen et al. (2017); Levi et al. (2016); Fischer et al. (2016); Medvedyeva et al. (2016); Hamazaki et al. (2019); Ren et al. (2020).
@@
Another class of entanglement phase transitions arises from the continuous monitoring of quantum systems. The stochastic Schrödinger equation typically governs these open quantum dynamics, and averaging over different realizations of the states leads to the well-known Lindblad master equation dynamics Paule (2018). Since entanglement entropy is a nonlinear function of the state, calculating it for each trajectory separately and then averaging reveals a phase transition. In contrast, any linear function of the state is featureless and does not exhibit such transitions. Several studies have explored these entanglement phase transitions in both interacting and non-interacting systems, considering clean as well as disordered systems with long-range and short-range hopping, using both Lindblad jump operators and quantum state diffusion approaches Alberton et al. (2021); Szyniszewski et al. (2023); Cao et al. (2019); Minato et al. (2022a); Müller et al. (2022); Yamamoto and Hamazaki (2023); Fuji and Ashida (2020); Ladewig et al. (2022). Continuously monitored open systems display competition between unitary evolution, which drives entanglement growth, and measurements, which tend to localize the wavefunction by collapsing it into eigenstates of the measurement operators. As a result, these systems undergo a measurement-induced phase transition as a function of measurement strength, which also has been reported recently in experiments Koh et al. (2023a). It has been shown that, in both clean and disordered free-fermionic systems, there is an entanglement phase transition from a critical to an area law phase as the strength of continuous measurements is tuned. This transition was predicted to fall into the Berezinskii-Kosterlitz-Thouless universality class Cao et al. (2019); Szyniszewski et al. (2023); Alberton et al. (2021), a result that was also supported by a theoretical study in which the dynamics of a continuously monitored free fermionic system were described by an effective non-Hermitian sine-Gordon model Buchhold et al. (2021). However, it is important to point out that the stability of the critical phase in a one-dimensional free-fermionic system under random projective measurements has been questioned recently, and it has been argued in the thermodynamic limit, for any finite measurement strength, only the area-law phase will survive Poboiko et al. (2023); Starchl et al. (2024a); Poboiko et al. (2025). Also in the case of Dirac fermions, the same conclusion is claimed recentlyMüller et al. (2025) after refining their previous resultsBuchhold et al. (2021). There are recent theoretical studies of a monitored chain of a system of size $L$ of Majorana fermions, which shows ($\ln L)^{2}$ behavior in the critical phase Fava et al. (2023), in contracts for noninteracting systems where the $U(1)$ symmetry is preserved, reported to show area-law phase for all measurement strength in presence of complex random hoppingFava et al. (2024). On the other hand, in the case of interacting systems, a transition from a volume law to an area law has been observed Fuji and Ashida (2020); Yamamoto and Hamazaki (2023); Xing et al. (2024). Furthermore, the role of measurements in localizing quantum states in nonunitary quantum circuits has also attracted significant interest Zabalo et al. (2022); Sierant and Turkeshi (2022); Oshima and Fuji (2023); Weinstein et al. (2023); Côté and Kourtis (2022); Block et al. (2022); Sang and Hsieh (2021); Lunt et al. (2021); Li et al. (2021); Fan et al. (2021); Zabalo et al. (2020); Turkeshi et al. (2020); Gullans and Huse (2020a, b); Bao et al. (2020); Choi et al. (2020); Tang and Zhu (2020); Jian et al. (2020). Some of these studies have also been extended in dimension $D>1$  Chahine and Buchhold (2024); Jin and Martin (2024); Poboiko et al. (2024) and a semi-classical limit of such transition has also been introduced recently Ruidas and Banerjee (2024). In contrast to disorder-induced localization observed in the Anderson-localized or MBL phase, the measurement-induced localization is governed by a completely different mechanism and driven by quantum measurements and measurement strength.
@@
In parallel, in the last few decades, periodically driven (Floquet) quantum systems have emerged as a powerful framework for studying non-equilibrium phases of matter Banerjee and Sengupta (2024); Sarkar et al. (2021, 2022); Nag et al. (2014); Agarwala et al. (2016); Luitz et al. (2018); Chakrabarty and Datta (2024a). Floquet systems are known to exhibit exotic phases, such as discrete-time crystals Khemani et al. (2016); Else et al. (2016); Zhang et al. (2017) and Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011); Kundu et al. (2014); Mukherjee et al. (2016). The fate of many-body localization (MBL) in the presence of a periodic drive has also been studied Lazarides et al. (2015); Ponte et al. (2015). While it has been demonstrated that Floquet systems can stabilize or even induce new phases, it remains an open question how continuous monitoring affects these driven systems and whether the introduction of a periodic drive can significantly alter the dynamics of a continuously monitored system, either enhancing or suppressing the effects of measurements. The effect of time-dependent Hamiltonian has been studied within the Lindblad master equation formalism Scopa et al. (2019); Ikeda et al. (2021), but the effect of periodic drive on a continuously monitored open quantum system, particularly in terms of trajectory-averaged entanglement entropy, remains relatively unexploredLi et al. (2019). In our study, we investigate such cases, and our main finding, based on our accessible system size numerics, is that periodically driven non-interacting systems tend to exhibit a signature of undergoing a BKT transition for reasonably high driving periods. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that if the BKT transition persists, then the critical measurement strength (transition point) increases with the time period. Moreover, the symmetric drive of the hopping amplitude about a zero mean is special; such a system always favors the area-law phase in the thermodynamic limit, independent of the frequency regime, in contrast to what has been observed in the case of the disorder-induced scenario Sarkar et al. (2021). We summarize our setup and main results based on the finite-sized numerical data in the form of a schematic in Fig. 1, which includes real data. We plot the maximum of the local effective central charge, which is a measure of the maximum growth rate of the steady state entanglement with sub-system size (defined later in the manuscript in Eq. (19)) with time-period $T$ for different measurement strength $\gamma$. Note that there is a growing belief based on recent analytical studies that in the no-drive scenario, any infinitesimal measurement strength is enough to stabilize the area law entanglement phase in the thermodynamic limit, but clear numerical evidence is still absent, specifically for the measurement protocols used in this manuscript. Moreover, to the best of our knowledge, no analytical prediction is available for driven systems till now. This work aims to do a detailed comparative numerical analysis between the driven system with a reasonably low-frequency regime (in the main text) and the no-drive/high-frequency regime (in the Appendix D) results for a finite-sized system.
@@
This manuscript is structured as follows. In Sec . II, we describe our system and protocols. Then, to build the intuitions about our systems, we focus on a $2\times 2$ toy model in Sec . III. Section IV is dedicated to the many-body systems, which consist of the analytical prediction for the high-frequency limit in sec. IV.1, as well as the detailed numerical results of the dependency of entanglement growth on driving periods in sec. IV.2. Section IV.3 concentrates on the fate of entanglement entropy growth on changing the asymmetry parameter $\epsilon$ of the square pulse drive as well as the fate of the sinusoidal driving protocol. Finally, we summarize our results in sec. V.
@@
Figure 1: Schematic of our protocol, where a free-fermionic lattice is continuously monitored, and the hopping amplitude $J(t)$ is periodically varied. Data is for the maximum of the local effective central charge vs time-period $T$ for different measurement strength $\gamma$ with $\epsilon=3J/4$.
@@ II Model and Methodology
where, $\hat{c}^{{\dagger}}_{i}$ and $\hat{c}_{i}$ are real-space fermionic creation and annihilation operators respectively. We study the effect of two different kinds of driving with the time period (frequency) $T$ ($\omega=1/T$): one is square pulse drive, where $J(t)=+J$ $(-J+\epsilon)$ for $20\leq t<T/$ 2 ($T/2\leq t<T$), with ϵitalic-$\epsilon\in$ $[0,2J]$ is an asymmetry parameter in the drive amplitude (see Fig. 1); the other is the sinusoidal drive where $J(t)=J\sin(2\pi t/T)$. Note that in the limit $\epsilon=0$, the mean of $J(t)$ is 00 over a time period, and identify it as a symmetric drive. We examine this limit separately in this manuscript. We primarily present the results for the square pulse in the subsequent sections of the manuscript; only at the end, in Sec. IV.3, some results for the sinusoidal drive are presented. For all our numerical calculations, we choose $J=1/2$. The time evolution of the state can be described using the stochastic Schrodinger equation as
@@
The above equation describes the continuous monitoring of the particle number $\hat{n}_{i}=\hat{c}_{i}^{{\dagger}}\hat{c}_{i}$ on each site, with measurement strength $\gamma$ Cao et al. (2019). $dW_{i}$s are taken from uncorrelated normal distributions having mean 00 and variance $\gamma dt$, which implies, $\overline{dW_{i,t}}=0$     and    $\overline{dW_{i,t}dW_{j,t^{\prime}}}=\gamma dt\delta_{i,j}\delta(t-t^{\prime})$. In our study, we use the quantum state diffusion protocol, while there are similar studies using the quantum jump protocol in clean as well as in disordered systems Alberton et al. (2021); Matsubara et al. (2025); Starchl et al. (2024b). For our model and protocols, the total number of particles $N=\sum_{i}\langle\hat{n}_{i}\rangle$ remain conserved in the entire time evolution process; hence, if we start with an initial state with a particular filling at $t=0$, for any time $t>0$, the filling fraction remains the same. In our work, we have fixed the filling fraction $N/L=1/2$ for all many-body calculations. Also, note that the last two terms in Eq. (LABEL:nlse) describe the contribution from the continuous measurement process and can be realized by homodyne detection in quantum optics Wiseman and Milburn (1993); Øksendal and Øksendal (2003).
@@
where, $\mathcal{M}=\sum_{i}[dW_{i}^{t}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt]\hat{n}_{i}$, and $\mathcal{N}$ is the normalization constant. Any pure Gaussian state can be represented as Cao et al. (2019),
@@
where $H(t)=J(t)\sum_{i=1}^{L}(\hat{c}_{i}^{{\dagger}}\hat{c}_{i+1}+h.c.)$ is the single particle Hamiltonian and $M_{ij}=\delta_{i,j}[dW_{i}^{t}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt]$. Under this non-Hermitian evolution, we need to properly normalize the wavefunction in each step, which can be done using the $QR$ decomposition of the $U$ matrix in each step as $U(t+dt)=QR$, and we set the new $U$ as $Q$. Since the columns of $Q$ remain orthonormal, this ensures that the quantum state remains normalized (see Appendix. E). Note that $U$ is an $L\crossproduct N$ matrix, follows $U^{{\dagger}}U=\mathbb{I}_{N\crossproduct N}$. On the other hand, the correlation matrix $D$ can be defined as $D_{ij}=\langle\hat{c}_{i}^{{\dagger}}\hat{c}_{j}\rangle=[UU^{{\dagger}}]_{ij}$ For our many-body calculations, we choose $dW_{i}$’s from different random realizations, compute the nonlinear function of states, e.g., entanglement entropy, using the correlation matrix for each realization separately, and finally average over different realizations (we ensure the convergence of our results by varying the number of realizations). On the other hand, the evolution of the average state will be governed by the Lindblad equation Cao et al. (2019).